A ball of mass 0.20 kg is fired with velocity 120 m/s into the barrel of a spring gun of mass 2.0 kg initially at rest on a frictionless surface. The ball sticks in the barrel at the point of maximum compression of the spring. No energy is lost to friction. What fraction of the ball's initial kinetic energy is stored in the spring?
Answer #1

At point of max compression the ball and gun have the same velocity (v)

Applying cons of momentum at max compression ..
0.2kg x 120m/s = (0.2+2.0)kg x v .. .. v = 24 / 2.2 = 10.90 m/s

Assuming perfect elastic collision with the spring (no energy loss) ..
Initial ball KE = final ball+gun KE + Energy in spring (Es)

Es = (½ x 0.2 x 120²) - (½ x 2.2 x 10.90²) =1309.30 J

Fraction of Ei in spring = Es / Ei = 1309.30 / (½ x 0.2 x 120²) .. .. ►0.90 J